# CasP: Improving Semi-Dense Feature Matching Pipeline Leveraging Cascaded Correspondence Priors for Guidance

### v1

@@ 1 Introduction
Figure 1: Matching Accuracy and Efficiency Comparisons. The runtime and AUC@20°$\degree$ accuracy are measured across two benchmarks and compared with LoFTR and ELoFTR (EL). The acceleration gains of ELoFTR diminish with increasing resolution since the matching stage occupies a significant portion of runtime. The novel cascaded matching pipeline is more efficient and robust than merely removing the dual-softmax (DS) operator.
@@ 3.2 Feature Interaction
Figure 4: Training-Inference Decoupling Cascaded Matching. During training, the top-$k$ priors selected by one-to-many matching include ground-truth correspondences for RSCA learning, and the DS operator is applied in both matching stages for supervision. During inference, one-to-many matching omits this step, whereas one-to-one matching applies partial softmax instead of the full version.
@@ 3.3 Cascaded Matching
Existing semi-dense methods [37, 33, 5, 7, 6] typically adopt the LoFTR [31] pipeline and apply a dual-softmax (DS) operator across both dimensions of the score matrix to filter out low-confidence matches. As evidenced by Fig. 1, this stage significantly increases the runtime, particularly for high-resolution inputs. ELoFTR [39] proposes a straightforward solution by omitting the DS operator and using the raw score matrix directly. While this strategy improves efficiency, it also diminishes robustness and generalizability, leading to reduced matching accuracy. To pursue a comprehensive solution, we propose a cascaded matching module that decouples training from inference, as illustrated in Fig. 4.
@@
We first construct the score matrix $S_{\nicefrac{{1}}{{16}}}$ from the correlations between $\mathbf{\widetilde{F}}_{\nicefrac{{1}}{{16}}}^{A}$ and $\mathbf{\widetilde{F}}_{\nicefrac{{1}}{{16}}}^{B}$. Our objective is to derive the top-$k$ correspondence priors $\pi_{\nicefrac{{1}}{{16}}}$ for each token in both views, which are defined as:
@@
Assuming one-to-one matches at the $\nicefrac{{1}}{{8}}$ scale, we set $k\geq 4$ because each token at the $\nicefrac{{1}}{{16}}$ scale may correspond to at most $(\nicefrac{{16}}{{8}})^{2}$ tokens in the cross-view feature map. During training, we apply a DS operator as a differentiable matching layer to $S_{\nicefrac{{1}}{{16}}}$, yielding distinctive feature representations and the confidence matrix $P_{\nicefrac{{1}}{{16}}}$ for supervision. In addition, we inject one-to-many ground-truth correspondences into $P_{\nicefrac{{1}}{{16}}}$ to accelerate convergence in the subsequent one-to-one matching. During inference, the DS operator is omitted because $S_{\nicefrac{{1}}{{16}}}$ alone suffices for the top-$k$ selection.
@@
@5°$\degree$ @10°$\degree$ @20°$\degree$ @5°$\degree$ @10°$\degree$ @20°$\degree$ on MD-1500 on SN-1500
@@
where $\chi_{\pi}(\cdot)$ denotes the indicator function and $\phi_{r}(\cdot)$ maps each position at the $\nicefrac{{1}}{{16}}$ scale to $r^{2}$ corresponding positions at the $\nicefrac{{1}}{{8}}$ scale. Coarse matches are filtered based on a predefined threshold $\theta$ over $P_{\nicefrac{{1}}{{8}}}$. The mutual-nearest-neighbor (MNN) criterion is then applied for one-to-one matching, forming $\mathcal{M}_{c}$. Note that Eq. (12) implies that each selected match $(i,j)$ must be derived from the correspondence priors on both sides, which can be formulated as follows:
@@ Results.
(0.25m,2°$\degree$) / (0.5m,5°$\degree$) / (1.0m,10°$\degree$) $\uparrow$
@@
(0.25m,2°$\degree$) / (0.5m,5°$\degree$) / (1.0m,10°$\degree$) $\uparrow$
@@ 4.4 Visual Localization
@5°$\degree$ @10°$\degree$ @20°$\degree$ @5°$\degree$ @10°$\degree$ @20°$\degree$ @5°$\degree$ @10°$\degree$ @20°$\degree$ @5°$\degree$ @10°$\degree$ @20°$\degree$
@@ 4.5 Understanding CasP
The ablation study primarily addresses two concerns: a) How does ELoFTR [39] perform when the DS operator is removed to speed up inference? b) As shown in Tab. 2, CasP achieves a more significant accuracy gain on SN-1500 than on MD-1500. How can this cross-domain generalization be further validated? To investigate these issues, we select two additional datasets, ETH3D[O] and ETH3D[I], from the zero-shot evaluation benchmark proposed by GIM [29]. These benchmarks represent real-world outdoor and indoor scenes in the ETH3D [28], respectively. The longer side of each image is resized to 1152 pixels, and standard RANSAC and MAGSAC++ [2] are employed as estimators. We draw the following conclusions from the results shown in Tab. 6 and  4.4: 1) Removing the DS operator is a trade-off that compromises accuracy, as the matching stage relies solely on descriptor similarity and ignores global confidence. 2) To examine the core design of our method, we replace the DS operator in ELoFTR with the cascaded matching module described in Sec. 3.3. Even the lite model performs comparably to or better than the original full model. 3) Building upon the novel matching pipeline, integrating additional advanced modules in our method further enhances accuracy. 4) Our pipeline requires fewer GMACs, delivers faster runtime, and uses less memory on MD-1500.

### v2

@@ 1 Introduction
Figure 1: Matching Accuracy and Efficiency Comparisons. The runtime and AUC@20$\degree$ accuracy are measured across two benchmarks and compared with LoFTR and ELoFTR (EL). The acceleration gains of ELoFTR diminish with increasing resolution since the matching stage occupies a significant portion of runtime. The novel cascaded matching pipeline is more efficient and robust than merely removing the dual-softmax (DS) operator.
@@ 3.2 Feature Interaction
[…]
@@ 3.3 Cascaded Matching
Existing semi-dense methods [38, 34, 5, 7, 6] typically follow the LoFTR [32] pipeline and apply a dual-softmax (DS) operator across both dimensions of the score matrix to filter out low-confidence matches. As evidenced by Fig. 1, this stage significantly increases the runtime, particularly for high-resolution inputs. ELoFTR [40] offers a simple solution by directly using the raw score matrix. While this strategy improves efficiency, it compromises robustness and generalizability. AdaMatcher [16] introduces a many-to-one assignment mechanism to address scale inconsistencies, but not for efficiency. To pursue a comprehensive solution, we propose a cascaded matching module, as illustrated in Fig. 4.
@@
[…]
@@
[…]
@@
@5$\degree$ @10$\degree$ @20$\degree$ @5$\degree$ @10$\degree$ @20$\degree$ on MD-1500 on SN-1500
@@
[…]
@@ Results.
(0.25m,2$\degree$) / (0.5m,5$\degree$) / (1.0m,10$\degree$) $\uparrow$
@@
(0.25m,2$\degree$) / (0.5m,5$\degree$) / (1.0m,10$\degree$) $\uparrow$
@@ 4.4 Visual Localization
@5$\degree$ @10$\degree$ @20$\degree$ @5$\degree$ @10$\degree$ @20$\degree$ @5$\degree$ @10$\degree$ @20$\degree$ @5$\degree$ @10$\degree$ @20$\degree$
@@ 4.5 Understanding CasP
The ablation study primarily addresses two concerns: a) How does ELoFTR [40] perform when the DS operator is removed to speed up inference? b) As shown in Tab. 2, CasP achieves a more significant accuracy gain on SN-1500 than on MD-1500. How can this cross-domain generalization be further validated? To investigate these issues, we select two additional datasets, ETH3D[O] and ETH3D[I], from the zero-shot evaluation benchmark proposed by GIM [30]. These benchmarks represent real-world outdoor and indoor scenes in the ETH3D [29], respectively. The longer side of each image is resized to 1152 pixels, and standard RANSAC and MAGSAC++ [2] are employed as estimators. We draw the following conclusions from the results shown in Tab. 6 and  7: 1) Removing the DS operator is a trade-off that compromises accuracy, as the matching stage relies solely on descriptor similarity and ignores global confidence. 2) To examine the core design of our method, we replace the DS operator in ELoFTR with the cascaded matching module described in Sec. 3.3. Even the lite model performs comparably to or better than the original full model. 3) Building upon the novel matching pipeline, integrating additional advanced modules in our method further enhances accuracy. 4) Our pipeline requires fewer GMACs, delivers faster runtime, and uses less memory on MD-1500.